Consumption of fast food is a topic of interest to researchers in the field of nutrition. An article reported that 1870 of those in a random sample of 6417 U.S. children indicated that on a typical day, they ate fast food. Estimate p, the proportion of children in the United States who eat fast food on a typical day. (Round your answer to four decimal places.)
Answer #1

Solution :

Given that,

n = 6417

x = 1870

Point estimate = sample proportion = = x / n = 1870/ 6417=0.291413

Point estimate =0.2914